# Hongzhi Zang1, Mingjie Wei6,2, Si Xu3, Yongji Wu5, Zhen Guo3, Yuanqing Wang4,3, Hao Lin3, &Liangzhi Shi1, Yuqing Xie1, Zhexuan Xu1, Zhihao Liu7,2, Kang Chen4,2, Wenhao Tang1, &Quanlu Zhang3, Weinan Zhang6, Chao Yu1,2,†,∗, Yu Wang1,†

### v1

@@ 3.1.3 Hybrid Allocation with Fine-grained Pipelining
This scheduling allows Simulator and Generation to run concurrently, reducing idle time while preserving correctness. In this way, our framework avoids the frequent offloading overhead of colocated allocation while also eliminating the GPU bubbles that occur in disaggregated allocation. Figure 3(c) provides a schematic illustration of the hybrid allocation mode with fine-grained pipelining ($k=2$).
3.1.4 Configuring GPU Allocation Modes
@@
Finally, fine-grained pipelining is controlled by the parameter rollout.pipeline_stage_num (corresponding to $k$ in Section 3.1.3). A value greater than 1 splits each simulator instance on a GPU into multiple pipeline stages, while a value of 1 disables pipelining.
3.2 Model Compatibility
Our framework is designed to be flexible and extensible, supporting both parameter-efficient training methods and different VLA model variants. In particular, we provide built-in support for Low-Rank Adaptation (LoRA) (Hu et al., 2021), as well as for the OpenVLA (Kim et al., 2024) and OpenVLA-OFT (Kim et al., 2025) models.
3.2.1 LoRA Support
LoRA (Hu et al., 2021) is a parameter-efficient fine-tuning method that adapts large pre-trained models to new tasks without updating all of their parameters. Instead of modifying the full weight matrices, LoRA introduces additional low-rank decomposition layers into linear transformations and trains only these small sets of parameters while keeping the original weights frozen. This greatly reduces the number of trainable parameters, lowering memory consumption and speeding up training while maintaining performance comparable to full fine-tuning. In our framework, users can flexibly choose whether to enable LoRA. If “is_lora” is enabled, all linear modules in VLA are designated as target modules by default, ensuring broad coverage of trainable parameters while preserving efficiency.
3.2.2 Model type
In RLinf-VLA, the argument “model_name” specifies the VLA model to be used, with current support for openvla” and “openvla_oft”. We briefly introduce these two models below.
OpenVLA
OpenVLA (Kim et al., 2024) is a VLA model containing about 7B parameters. Despite being smaller than some closed-source counterparts such as RT-2-X (55B) (O’Neill et al., 2024), OpenVLA demonstrates superior performance on a range of robotic manipulation benchmarks, achieving higher success rates while remaining computationally efficient. Our framework supports training and evaluation of OpenVLA, with extensive experiments conducted in the ManiSkill environment.
OpenVLA-OFT
OpenVLA-OFT (Optimized Fine-Tuning) (Kim et al., 2025) extends OpenVLA by focusing on fine-tuning efficiency, inference speed, and practical deployability in robotics. Instead of relying on discrete action tokens and complex objectives, OpenVLA-OFT represents actions in a continuous space and optimizes them using a simple L1 regression loss. This design reduces training overhead and accelerates convergence. Moreover, OpenVLA-OFT incorporates parallel decoding together with action chunking, which significantly improves inference throughput and enables real-time control at higher frequencies. These advances yield strong performance gains in both the LIBERO benchmark and real-world tasks, while also achieving over an order-of-magnitude speedup in action generation. Our framework supports training OpenVLA-OFT with a discretized action mode and provides experimental results on ManiSkill and LIBERO environments.
3.2.3 Unified Interface
3.3 Multiple Simulators Support
Simulators play a crucial role in reinforcement learning for robotics, providing controllable and efficient environments in which agents can interact, learn, and improve without the high costs and risks of physical hardware. Compared to real-world experiments, simulators offer several key advantages:
  • Scalability of data collection: By leveraging parallel computation, simulators can generate a large volume of interaction data (often at least thousands of frames per second), far exceeding what is feasible with physical robots.
  • Task diversity: Simulators provide access to a wide range of robotic tasks, including table-top manipulation, drawing/cleaning, and dexterous manipulation. This diversity is crucial for improving and evaluating the generalization capabilities of VLA models.
  • Safety and controllability: Agents can explore potentially dangerous or failure-prone behaviors without risking damage to hardware.
  • Fast environment reset and reduced human labor: Simulators enable instant and reproducible environment resets, greatly accelerating the training cycle and eliminating the need for manual intervention during repeated experiments.
Given these advantages, our framework focuses on implementing robotics RL in simulation. In this work, we select two representative platforms, ManiSkill and LIBERO, as our primary simulators. ManiSkill excels in physics-based manipulation tasks with diverse object assets, while LIBERO emphasizes instruction-conditioned tasks that test generalization and high-level reasoning for VLA models. Together, these simulators provide complementary capabilities, forming a solid foundation for training and evaluating our framework.
3.3.1 Unified Interface
Our framework supports vectorized environments with a unified interface across different simulators. For example, we leverage ManiSkill’s GPU-parallelized vectorized environments and LIBERO’s CPU-parallelized vector environment wrapper. On top of these backends, we define a consistent interface to seamlessly integrate with our training framework.
Core Functions
We implement standard Gym-style APIs, including “reset” and “step”. The “step” function additionally supports an “auto_reset” option: when enabled, any sub-environment that terminates or is truncated will be automatically reset, thereby improving sample efficiency by avoiding idle environments. Following the ManiSkill convention, we also support the “ignore_terminations” option. When enabled, termination signals are ignored and only truncation signals are respected, meaning that an episode ends only when the maximum episode length is reached. This feature allows us to flexibly support different implementation variants discussed in Section 4.2.1 and Section 4.2.2.
Utility Functions
3.3.2 Tasks
We implement a variety of tasks based on the two simulators in our framework.
ManiSkill
Following RL4VLA (Liu et al., 2025), we construct the environment “PutOnPlateInScene25Main-v3”, which consists of 25 pick-and-place tasks with variations in objects and receptacles. We also adopt the out-of-distribution evaluation settings proposed in RL4VLA. By replicating their results, we further confirm the generalizability of VLA-RL models under these task variations.
LIBERO
We leverage the public benchmark task groups provided by LIBERO (Liu et al., 2023b), including LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, LIBERO-10, and LIBERO-90. Specifically:
  • LIBERO-Spatial: evaluates the model’s ability to reason about relative object positions.
  • LIBERO-Object: tests generalization across different object types.
  • LIBERO-Goal: examines the model’s capability to adapt to diverse task goals.
  • LIBERO-10: a long-horizon benchmark consisting of 10 tasks.
  • LIBERO-90: a long-horizon benchmark covering 90 tasks.
Most existing VLA+RL studies train and evaluate models within a single task group. In contrast, we train our model on the combined benchmark, LIBERO-130, which comprises 130 tasks specifically constructed to evaluate RLinf-VLA’s capability to learn under large-scale task settings, and we assess its performance across these task groups.
3.4 Multi-Algorithm Support
Our framework provides support for multiple reinforcement learning algorithms, with an initial focus on Proximal Policy Optimization (PPO) (Schulman et al., 2017) and Group Relative Policy Optimization (GRPO) (Shao et al., 2024; Guo et al., 2025). Both are representative on-policy algorithms that have been widely adopted in robotics research. PPO has established itself as a robust and general-purpose baseline for continuous control, while GRPO offers a simplified alternative by removing the need for value function estimation. Together, they form a solid foundation for training and evaluating VLA models in simulators.
3.4.1 PPO
PPO is one of the most widely adopted reinforcement learning algorithms in robotics. PPO enhances training stability by constraining policy updates within a trust region, thereby preventing overly large changes that could destabilize learning. This is achieved through a clipped surrogate objective, which balances exploration and exploitation while maintaining sample efficiency. Due to its robustness and simplicity, PPO has become a standard baseline in robotics RL and serves as the foundation of our framework.
In PPO, the advantage function is commonly estimated using Generalized Advantage Estimation (GAE) (Schulman et al., 2015):
The PPO optimization objective is defined as:
where
$\pi_{\theta_{\text{old}}}$ denotes the rollout policy, and $\epsilon$ is the clipping parameter. [^2] The definition of action $a_{t}$ differs across LLM and robotics RL settings. Here we treat it as a generalized action, with the distinctions discussed in Section 4.1.2. The same applies to GRPO.
3.4.2 GRPO
GRPO is a recent variant of policy optimization designed to simplify reinforcement learning pipelines. Unlike PPO, which requires training both a policy model and a value model, GRPO eliminates the need for an explicit value function by leveraging group-based relative comparisons of trajectories. This design reduces the overall model complexity and avoids potential inaccuracies from value estimation. As a result, GRPO provides a lightweight yet effective alternative to PPO, making it especially attractive for large-scale VLA training where efficiency and simplicity are important. We begin with a brief overview of GRPO as applied in LLM reasoning models, and present its specific design for robotics in Section 4.1.2. In GRPO, the important ratio can be defined as:
where $i$ denotes $i$-th trajectory. But the advantage can be defined as:
4 Algorithm Design
4.1 Advantage and Log-Probability
4.1.1 Advantage with Action Chunks
Figure 4: Illustration of different logprob levels.
Action Chunk
An action chunk refers to having the policy predict a short sequence of future actions rather than a single-step action at each timestep. This approach has been shown to better handle non-Markovianity in offline data, such as the non-Markovian behaviors often present in human demonstrations (Zhao et al., 2023). The technique is widely adopted in imitation learning. To integrate action chunks into reinforcement learning, we must define how advantages are associated with actions inside a chunk. Consider the $t$-th chunk $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$, where $C$ denotes the chunk length. Each atomic action $a_{t,j}$ produces an immediate reward $r_{t,j}$. Two formulations are commonly used:
  1. 1.
    Chunk-level: The entire chunk $c_{t}$ is treated as a macro-action, with the reward defined as the sum of per-step rewards, i.e., $r_{t}=\sum_{j=1}^{C}r_{t,j}$.
  2. 2.
    Action-level: Each $a_{t,j}$ is treated as an individual action, with its own reward and advantage.
These two formulations lead to different behaviors under GAE-based estimation.
Our framework supports both options. Users can control the advantage formulation via the parameter reward_type, which can be set to either chunk_level or action_level.
4.1.2 Log-Probability Granularity
In algorithms such as PPO and GRPO, computing the log-probabilities of actions under the current policy is essential for estimating the importance sampling ratios $\rho(\theta)$ in Equations (8) and (9). However, the precise definition of an “action” depends on the task domain.
In robotics RL, probabilities are typically computed either at the action level or the chunk level, depending on whether each environment step is considered atomic or grouped into a chunk. In contrast, in LLM-based reasoning tasks, policies generate sequences of tokens, where the natural atomic unit is a token. In this case, action probabilities are computed at the token level, by multiplying (or summing in log-space) the probabilities of individual tokens.
Thus, our framework supports three granularity choices for log-probability computation: action_level, chunk_level, and token_level, as shown in Figure 4. The selection is controlled by the parameter logprob_type, and each option leads to different optimization dynamics. Formally, we define a trajectory as $\tau=(c_{0},c_{1},\ldots,c_{T})$, where each chunk $c_{t}$ contains $C$ atomic actions, $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$, and each atomic action $a_{t,i}$ consists of $M$ tokens, $a_{t,i}=(d_{t,i,1},d_{t,i,2},\ldots,d_{t,i,M})$. The policy probabilities decompose as:
  • Chunk-level: $\pi(c_{t}|o_{t})=\prod_{i=1}^{C}\pi(a_{t,i}|o_{t},a_{t,:i-1})$,
  • Action-level: $\pi(a_{t,i}|o_{t},a_{t,:i-1})=\prod_{j=1}^{M}\pi(d_{t,i,j}|o_{t},d_{t,i,:j-1})$,
  • Token-level: $\pi(d_{t,i,j}|o_{t},d_{t,i,:j-1})$.
4.1.3 Supported Granularity Combinations for Advantage and Log-Probability
Table 1 summarizes the supported combinations of advantage and log-probability granularities. In general, an advantage type is compatible with all log-probability types whose granularity is no finer than its own. When the log-probability is defined at a finer granularity, we apply a broadcasting mechanism to the advantage values. For instance, if the advantage is defined at the chunk level while the log-probability is computed at the token level, the same chunk-level advantage is assigned to each token within the chunk.
Table 1: Supported combinations of advantage and log-probability granularities.

### v2

@@ IV-A3 Hybrid Allocation with Fine-grained Pipelining
This scheduling allows Simulator and Generation to run concurrently, reducing idle time while preserving correctness. In this way, our framework avoids the frequent offloading overhead of collocated allocation while also eliminating the GPU bubbles that occur in disaggregated allocation. Figure 4 provides a schematic illustration of the hybrid allocation mode with fine-grained pipelining ($k=2$).
IV-B A Unified Interface
To conveniently reuse existing simulators, models, and algorithms, RLinffff-VLA provides a unified interface designed for high extensibility and scalability.
IV-B1 Multiple Simulators Support.
RLinff-VLA supports multiple robotic simulators, including ManiSkill [39], LIBERO [21], and RoboTwin [8]. To facilitate this, we provide a unified interface that standardizes environment interactions across different backends, such as ManiSkill and RoboTwin’s GPU-parallelized environments and LIBERO’s CPU-parallelized wrappers. This interface includes standard Gym-style APIs with built-in support for action chunking and flexible episode termination. Furthermore, we define a unified interface for observation and action to ensure seamless compatibility with various algorithms and models.
IV-B2 Multiple Models Support
Benefiting from a unified data processing pipeline across simulators, our framework facilitates the seamless integration of existing models by requiring only a standardized interaction interface. We support diverse VLA architectures, exemplified in this study by OpenVLA [18] and OpenVLA-OFT [17], which represent configurations with single-step and multi-step action chunking, respectively. To enable efficient fine-tuning, we integrate Low-Rank Adaptation (LoRA) [13].
IV-B3 Multiple Algorithms Support
Our framework provides support for multiple reinforcement learning algorithms, with an initial focus on Proximal Policy Optimization (PPO) [33] and Group Relative Policy Optimization (GRPO) [35, 11]. The core functions for on-policy RL algorithms are the advantage function and the loss function. In our framework, we can configure the algorithms by specifying these two functions only, without redundant code.
IV-C Design Choices for Algorithms
RL algorithms for large models involve more complex design choices than those for smaller models. For VLAs, we adopt methodologies from both LLMs and robotics. In Section IV-C1, we introduce general features applicable to all algorithms. Subsequently, in Section IV-C2 and Section IV-C3, we detail specific design choices for PPO and GRPO, respectively. Consistent with our codebase style, all these features can be easily configured by changing values in the configuration file.
IV-C1 Multi-Granularity Calculation Support
Our framework supports advantage estimation and log-probability computation at multiple levels of granularity, enabling seamless adaptation to different algorithmic requirements.
For advantage estimation, we provide two formulations for assigning advantages to action chunks $c_{t}=(a_{t,1},\dots,a_{t,C})$ [^1] Here, $t$ denotes the index of the chunk, and $C$ denotes the chunk size.. In the chunk-level formulation, the entire chunk is treated as a single macro-action and is assigned a summed reward and a shared advantage. In contrast, the action-level formulation assigns individual rewards and advantages to each atomic action $a_{t,i}$ within the chunk.
Similarly, for log-probability computation, we support three corresponding granularities:
Chunk-level: $\;\pi(c_{t}|o_{t})=\textstyle\prod_{i=1}^{C}\pi(a_{t,i}|o_{t},a_{t,{:}i-1}),$
Action-level: $\;\pi(a_{t,i}|o_{t},a_{t,{:}i-1})=\textstyle\prod_{j=1}^{M}\pi(d_{t,i,j}|o_{t},d_{t,i,{:}j-1}),$
Token-level: $\;\pi(d_{t,i,j}|o_{t},d_{t,i,{:}j-1}),$
where $d_{t,i,j}$ denotes the $j$-th token of the $i$-th atomic action.
(a) Fixed Length
(b) Partial Reset
(c) Valid Action Mask
Figure 6: Illustration of the three rollout modes.
IV-C2 Design Choices for PPO
Scaling PPO to VLA models introduces several design challenges. We summarize the key design decisions adopted in our framework below.
Partial Reset Support During rollouts, a sub-environment typically terminates either upon reaching the maximum episode length or upon successfully completing a task. We consider two standard strategies for handling such terminations. In the Fixed Episode Length mode, all sub-environments are reset simultaneously only after reaching the maximum episode length. In contrast, the Partial Reset mode resets each sub-environment immediately upon termination, independent of others. Figure 6(a) and Figure 6(b) illustrate the differences between these two modes.
Critic Design Adapting PPO to VLA models places stringent requirements on critic design. Maintaining a separate critic network is computationally prohibitive and complicates GPU resource management due to the large scale of VLA models. We therefore adopt a parameter-sharing strategy between the actor and critic. Following RL4VLA [22], we attach a lightweight value head to the language model component of the VLA for efficient state value estimation.
Value for Action Chunks As discussed in Section IV-C1, the two advantage estimation formulations naturally induce two corresponding value estimation strategies for an action chunk $c_{t}=(a_{t,1},a_{t,2},\ldots,a_{t,C})$. In the chunk-level formulation, the critic estimates a single scalar value for the entire chunk, treating it as a macro-action, i.e., $V:\mathcal{S}\rightarrow\mathbb{R}$. In contrast, the action-level formulation produces a $C$-dimensional value vector, providing an individual value estimate for each atomic action $a_{t,i}$ in the chunk, i.e., $V:\mathcal{S}\rightarrow\mathbb{R}^{C}$. Empirically, we find that the action-level formulation consistently leads to better performance.
IV-C3 Design Choices for GRPO
Transferring GRPO to embodied tasks introduces several non-trivial challenges. To address these, we adopt the following design choices.
Valid Action Mask. While rollouts are executed for a fixed episode duration, tasks often terminate early. This presents two potential strategies for policy optimization: (1) using the full trajectory regardless of task completion timing, or (2) considering only timesteps prior to task completion. Our framework supports both strategies, referring to the latter as the Valid Action Mask setting (Figure 6(c)). Empirically, we find that masking out post-completion steps generally improves policy performance in the GRPO setting.
Loss Normalization by Trajectory Length. To ensure that trajectories of different lengths contribute comparably to optimization, we normalize the policy loss by the number of valid timesteps in the Valid Action Mask setting. Specifically, for a trajectory $\tau_{i}$ with $T_{i}^{\text{succ}}$ valid timesteps, the contribution of each timestep to the objective is scaled by $1/T_{i}^{\text{succ}}$. This prevents longer trajectories from dominating the gradient and promotes balanced learning across successful and partially completed trajectories.
Success Rate Filter. Inspired by the dynamic sampling strategy in DAPO [42], we introduce a success-rate filter for GRPO. This filter discards trajectory groups where all trajectories either succeed or fail, as computing non-zero advantages requires a mix of successful and failed outcomes. In practice, this mechanism accelerates convergence and consistently improves policy performance.
(a) OpenVLA on ManiSkill.
(b) OpenVLA-OFT on ManiSkill.
(c) OpenVLA-OFT on LIBERO.
(d) OpenVLA-OFT on RoboTwin.
Figure 7: Training curves on different benchmarks. The x-axis shows the number of training epochs, and the y-axis indicates the corresponding success rate. Light-colored lines represent the raw data, while the dark-colored curves are smoothed using a Gaussian filter with $\sigma=1$.
V Experiment Results
In this section, we systematically investigate three key questions concerning the effectiveness of the proposed RLinf-VLA framework:
(1) Is RLinf-VLA high-performance? We evaluate RLinf-VLA on three representative testbeds, LIBERO, ManiSkill and RoboTwin. The results demonstrate that RLinf-VLA achieves approximately 20–85% improvement, highlighting its strong capability to support large-scale multi-task learning.
(2) Is RLinf-VLA high-efficiency? We benchmark the framework across both GPU-parallelized and CPU-parallelized simulators, and observe that the optimal configuration improves training throughput, with speedups of up to $1.88\times$. This finding underscores the necessity of supporting diverse allocation modes. Notably, RLinf-VLA achieves up to $2.27\times$ speedup over existing frameworks.
(3) What are the actionable practices for applying PPO and GRPO to VLA training? Through extensive ablation studies, we identify the key factors that govern training performance, offering practical guidelines for effectively deploying RL in VLA settings.
TABLE I: Evaluation results across three simulation benchmarks. Values denote success rates (%).
(a) Evaluation on ManiSkill.
Method In-Distribution OOD Avg.
OpenVLA (Base) 53.91 39.10
OpenVLA (RLinf-GRPO) 84.38 75.15
OpenVLA (RLinf-PPO) 96.09 81.93
OpenVLA-OFT (Base) 28.13 18.29
OpenVLA-OFT (RLinf-GRPO) 94.14 60.64
OpenVLA-OFT (RLinf-PPO) 97.66 77.05
(b) Evaluation on LIBERO.
OpenVLA-OFT Object Spatial Goal Long 90 Avg.
Base 50.20 51.61 49.40 11.90 42.67 42.09
RLinf-GRPO 99.67 98.93 98.32 93.55 98.12 98.11
(c) Evaluation on RoboTwin.
OpenVLA-OFT Cup Hammer Bottles Can Pot Hand Avg.
Base 75.78 10.15 20.31 9.37 3.13 28.13 24.48
RLinf-GRPO 94.53 96.09 92.96 83.59 70.31 70.31 84.63
V-A Policy Performance
In this section, we evaluate the effectiveness of RLinf-VLA policies across multiple benchmarks, focusing on task success rates and generalization performance.
Experimemt Setup. We evaluate RLinf-VLA on three benchmarks: ManiSkill, LIBERO, and RoboTwin, with full settings in Appendix III.A.
For ManiSkill, we train on 25 pick-and-place tasks with object and receptacle variations. We follow the out-of-distribution evaluation protocol of RL4VLA [22], measuring generalization in vision, language, and action. Each sub-setting is evaluated with 256 random episodes. For LIBERO, we use the public task groups [21], including LIBERO-Spatial, Object, Goal, Long, and 90, for training and evaluation. Instead of training within a single group, we train one unified model on the combined 130 tasks over 5 groups, testing large-scale multitask learning. Performance is reported across groups, averaged over 50 episodes per task with three random seeds. For RoboTwin [8], we select six tasks: “beat block hammer”, “move can pot”, “place empty cup”, “pick dual bottles”, “lift pot”, and “handover block”. Training uses 1,000 fixed randomized scene seeds per task, and evaluation samples 128 unseen seeds to assess out-of-distribution generalization.
Training Results. Figure 7 presents the training curves for each setup. Figure 7(a) and Figure 7(b) report the training performance on ManiSkill, utilizing OpenVLA and OpenVLA-OFT models trained via PPO and GRPO. Across all settings, RL delivers substantial performance gains, improving success rates by 45%–70% over the baseline. Notably, PPO consistently outperforms GRPO and exhibits greater stability for both OpenVLA and OpenVLA-OFT in the ManiSkill setting.
Figure 7(c) illustrates the training curve of OpenVLA-OFT with the GRPO algorithm on 130 tasks of LIBERO. Results demonstrate that the success rate improves substantially from approximately $73\%$ to $98\%$, yielding an overall performance gain of about $30\%$. These findings underscore the effectiveness of the GRPO design in RLinf-VLA for enhancing multi-task training.
Figure 7(d) depicts our training curves for OpenVLA-OFT on RoboTwin. The training process remains stable with minimal fluctuations. Moreover, the model converges to high performance even on tasks with extremely low initial success rates (as low as 3%).
(a) Throughput for different settings.
(b) Latency breakdown.
Figure 8: Evaluation of system efficiency on ManiSkill, LIBERO, and RoboTwin, comparing OpenVLA and OpenVLA-OFT. Throughput (total environment frames per second) improves with increasing pipeline stages (“pipe”).
Evaluation Results. Table I summarizes RLinf-VLA results across three benchmarks, with ManiSkill detailed in Table I(a). “OpenVLA (Base)” and “OpenVLA-OFT (Base)” denote the RL base models for OpenVLA and OpenVLA-OFT.
Direct OOD comparison between OpenVLA and OpenVLA-OFT is not strictly fair because their base performance differs. OpenVLA improves from $39.10\%$ to $81.93\%$ after RL, while OpenVLA-OFT rises from $18.29\%$ to $77.05\%$. Although starting weaker, OpenVLA-OFT shows slightly larger relative gains, indicating that base model choice strongly affects final generalization.
Table I(b) reports OpenVLA-OFT trained on 130 tasks with RLinf-VLA over five LIBERO groups. LIBERO-Object reaches $3\%99$$9\%$ success, with an overall average of $98.11\%$ and a mean improvement of $56.02\%$. Notably, we use a single unified model across all tasks, unlike standard RL-for-VLA methods that train separate models per group, demonstrating RLinf-VLA’s support for large-scale multi-task RL.
OOD results on RoboTwin are shown in Table I(c). RLinf-VLA achieves an average success rate of $84.63\%$, over $60\%$ improvement, showing strong generalization on complex tasks such as long-horizon bimanual manipulation.
V-B System Efficiency
In this section, we evaluate the efficiency of our system by comparing different GPU allocation strategies. Our results show that the optimal strategy varies across simulators and models, highlighting the importance of flexible GPU allocation, which is a core feature of RLinf-VLA.
Experiment Setup. We summarize key settings here, with full details in Appendix III.B. We evaluate representative tasks from three benchmarks. For ManiSkill, we use “PutCarrotOnPlateInScene-v2” with 256 parallel environments. For LIBERO, we adopt the LIBERO-Long suite, and for RoboTwin, the “place empty cup” task. For LIBERO and RoboTwin, parallel environments scale linearly with nodes (64, 128, and 256 for 1, 2, and 4 nodes).
We consider three GPU allocation strategies (Section IV-A). Collocated shares GPUs across components, Disaggregated separates training from rollout, and Hybrid pipelines simulation by splitting instances into stages. For comparisons, ManiSkill lacks multi-GPU baselines, so we compare Collocated and Hybrid (pipelined) modes against a naive Disaggregated baseline. For LIBERO and RoboTwin, we benchmark against SimpleVLA-RL [20] (collocated only) and evaluate our Collocated and Hybrid modes; the disaggregated mode is omitted since hybrid consistently performs better when GPUs are underutilized.
Results. Figure 8 reports the end-to-end throughput of RLinf-VLA and baselines for VLA+RL training under different cluster sizes and placement strategies.
[…]
For OpenVLA in ManiSkill and OpenVLA-OFT in ManiSkill, as GPU count increases, the hybrid (pipe=1) mode outperforms the collocated mode. Since ManiSkill is GPU-parallelized, rollout throughput scales with parallel environments, making simulator resource allocation critical. To improve utilization, we enable offloading in the collocated mode (Section IV-A1), but communication overhead grows with cluster size. In contrast, hybrid mode splits GPUs evenly between components, reducing interference.
For OpenVLA-OFT in LIBERO and ManiSkill, the behavior differs: collocated and hybrid (pipe=1) modes outperform disaggregated and hybrid (pipe=2). First, OpenVLA-OFT generates action chunks while the simulator executes them sequentially, shifting the generation-to-execution ratio from 1:1 to about 1:15. Second, LIBERO’s CPU-parallelized simulator becomes the main bottleneck. This imbalance diminishes the benefit of pipelining, effectively reverting execution to a near-sequential process.
[…]
Summary. Different simulators demand different allocation strategies. RLinf-VLA’s flexible multi-mode support enables consistently high efficiency across diverse execution regimes.
(a) succ. in ManiSkill
(b) loss in ManiSkill
(c) succ. in LIBERO
Figure 9: For OpenVLA-OFT, action-level value estimation consistently outperforms chunk-level estimation across different tasks.
V-C Ablation Study
In this section, we present ablation studies under different setups and summarize key insights for PPO and GRPO training.
(1) Tips for PPO.
(a) Action-level value estimation outperforms chunk-level estimation for PPO with action chunks. Figure 9 shows the PPO training curves comparing action-level and chunk-level value estimation as described in Section IV-C2 on the ManiSkill “PutOnPlateInScene25Main” task using the OpenVLA-OFT model. Action-level estimation consistently yields higher success rates and lower value loss, demonstrating more effective learning and faster policy improvement. The performance divergence between different levels of value estimation remains consistent across diverse tasks, as corroborated by similar results in the LIBERO-Goal benchmark (Figure 9(c)).
(a) succ. in OpenVLA
(b) succ. in OpenVLA-OFT
Figure 10: Partial reset consistently outperforms Fixed Episode Length mode in PPO.
(b) Partial reset substantially improves sample efficiency. Figure 10 presents the PPO training curves for the Partial Reset and Fixed Episode Length rollout modes discussed in Section IV-C2. Since the optimization objective is to achieve success at least once per rollout (“success_once”), Partial Reset leads to a significantly higher success rate. For a given number of training steps, the success rate under Partial Reset consistently exceeds that of the Fixed Episode Length mode. This trend is observed regardless of the model type (OpenVLA or OpenVLA-OFT).
Figure 11: Ablation study on trajectory length normalization. Success rate in LIBERO-Goal with OpenVLA-OFT.
(2) Tips for GRPO.
(a) Trajectory length normalization in GRPO. As discussed in Section IV-C3, normalizing the loss by trajectory length aims to reduce bias when episodes vary in length. Figure 11 shows that incorporating trajectory length normalization (“w/ Norm”) can lead to substantially higher performance compared with the unnormalized setting (“w/o Norm”).
(b) Valid action mask in GRPO. We also study the effect of valid action masking, introduced in Section IV-C3. Since the optimization objective is defined with respect to “success_once”, excluding actions after reaching the success state improves sample efficiency and avoids redundant updates. Moreover, applying the valid action mask naturally results in shorter trajectories, which further benefits trajectory length normalization. Figure 12(a) shows results on LIBERO-Goal. Comparing the “w/o Mask, w/o Norm” and “w/ Mask, w/o Norm” curves, the setting with a valid action mask achieves consistently better performance. The “w/ Mask, w/ Norm” curve demonstrates that combining the mask with trajectory length normalization provides an additional improvement. However, the effect of these techniques is task-dependent. In the ManiSkill setting, we do not observe clear benefits from either valid action masking or trajectory length normalization.
(a) succ. in LIBERO-Goal
(b) succ. in ManiSkill
Figure 12: Ablation studies on valid action mask in GRPO with OpenVLA-OFT.
(c) Success rate filtering can improve training stability in some settings of GRPO. As discussed in Section IV-C3, we implement a success rate filter for GRPO that discards groups in which all trajectories have identical cumulative rewards. This mechanism can improve the stability of GRPO training. For instance, in the OpenVLA ManiSkill setting, training without the filter (“w/o Filter”) exhibits a clear collapse around step 400, whereas enabling the filter (“w/ Filter”) largely alleviates this issue. However, the benefit of the filter is not universal: in the OpenVLA-OFT ManiSkill setting and the OpenVLA-OFT LIBERO-Goal setting, its effectiveness is much less pronounced.
VI Conclusion
In this work, we introduced RLinf-VLA, a unified and efficient framework for reinforcement learning-based training of Vision-Language-Action models. RLinf-VLA integrates multiple simulators, algorithms, and VLA architectures, while providing flexible execution modes and system-level optimizations that significantly improve training efficiency. Extensive experiments demonstrate that models trained with RLinf-VLA achieve approximately 20–85% improvement on a wide range of simulated tasks. Moreover, our study distills actionable practices for both PPO and GRPO, guiding future research in RL-based VLA training. By open-sourcing RLinf-VLA with ongoing maintenance, we provide the community with a foundation to accelerate, standardize, and scale research in embodied intelligence.
Appendix for RLinf-VLA
VII Theoretical Background
VII-A Definitions of RL
Definition 1 (Return).
$R_{t}=\sum_{k=0}^{\infty}\gamma^{k}r(s_{t+k},a_{t+k}).$ (2)
Definition 2 (Value Function).
The value function $V_{\pi}(s)$ is the expected return when starting from state $s$ and following policy $\pi$:
$V_{\pi}(s)=\mathbb{E}_{\pi}[R_{t}\mid s_{t}=s],$ (3)
where $t$ denotes an arbitrary timestep at which the agent is in state $s$.
Definition 3 (Action-Value Function).
The action-value function $Q_{\pi}(s,a)$ is the expected return when executing action $a$ in state $s$ and thereafter following policy $\pi$:
$Q_{\pi}(s,a)=\mathbb{E}_{\pi}[R_{t}\mid s_{t}=s,a_{t}=a],$ (4)
where $t$ denotes an arbitrary timestep at which the agent is in state $s$ and takes action $a$.
Definition 4 (Advantage Function).
$A_{\pi}(s,a)=Q_{\pi}(s,a)-V_{\pi}(s).$ (5)
VII-B Introduction of PPO
PPO [33] is one of the most widely adopted reinforcement learning algorithms in robotics. PPO enhances training stability by constraining policy updates within a trust region, thereby preventing overly large changes that could destabilize learning. This is achieved through a clipped surrogate objective, which balances exploration and exploitation while maintaining sample efficiency. Due to its robustness and simplicity, PPO has become a standard baseline in robotics RL and serves as the foundation of our framework.
In PPO, the advantage function is commonly estimated using Generalized Advantage Estimation (GAE) [32]:
The PPO optimization objective is defined as:
where
$\pi_{\theta_{\text{old}}}$ denotes the rollout policy, and $\epsilon$ is the clipping parameter. [^2] The definition of action $a_{t}$ differs across LLM and robotics RL settings. Here we treat it as a generalized action. The same applies to GRPO.
VII-C Introduction of GRPO
GRPO [35] is a recent variant of policy optimization designed to simplify reinforcement learning pipelines. Unlike PPO, which requires training both a policy model and a value model, GRPO eliminates the need for an explicit value function by leveraging group-based relative comparisons of trajectories. This design reduces the overall model complexity and avoids potential inaccuracies from value estimation. As a result, GRPO provides a lightweight yet effective alternative to PPO, making it especially attractive for large-scale VLA training where efficiency and simplicity are important. In GRPO, the important ratio can be defined as:
where $i$ denotes $i$-th trajectory. But the advantage can be defined as:
VIII Implementation Details of RLinf-VLA
VIII-A The Unified Interface
Core Functions
We implement standard Gym-style APIs, including “reset” and “step”. The “step” function additionally supports an “auto_reset” option: when enabled, any sub-environment that terminates or is truncated will be automatically reset, thereby improving sample efficiency by avoiding idle environments. Following the ManiSkill [39] convention, we also support the “ignore_terminations” option. When enabled, termination signals are ignored and only truncation signals are respected, meaning that an episode ends only when the maximum episode length is reached. This feature allows us to flexibly support different implementation variants, such as “Partial Reset” and “Valid Action Mask”.
Utility Functions
VIII-B GPU Allocation Modes
Configuring GPU Allocation Modes
@@ VIII-B GPU Allocation Modes
Finally, fine-grained pipelining is controlled by the parameter rollout.pipeline_stage_num. A value greater than 1 splits each simulator instance on a GPU into multiple pipeline stages, while a value of 1 disables pipelining.
VIII-C Details for Design Choices
VIII-C1 Details for Multi-Granularity Support
In the main text, we refer to three levels of granularity: token-level, action-level, and chunk-level. Figure 13 provides an illustrative example of these levels. An action chunk consists of multiple atomic actions, which are the actual control signals executed by the simulator at each step. Each atomic action, in turn, is composed of multiple tokens, where each token corresponds to a single action dimension.
Figure 13: Illustration of different log-probability granularities.
The complete set of supported combinations between advantage computation granularity and log-probability granularity is summarized in Table II.
TABLE II: Supported combinations of advantage and log-probability granularities.